OpenStudy (priyar):

1^2 +3^2 +5^2 +.....+25^2=?

OpenStudy (priyar):

i did summation of (2n + 1)^2

OpenStudy (priyar):

from 0 to 25 and substituted the values...

OpenStudy (faiqraees):

shouldn't it be (2n-1)^2

OpenStudy (priyar):

*0 to 12

OpenStudy (priyar):

it could be either way i guess but if its 2n -1 we must start from 1...

OpenStudy (faiqraees):

oh you were starting from 0 then its also correct

OpenStudy (jiteshmeghwal9):

sum of squares of first n odd natural numbers\[\frac{n(2n+1)(2n-1)}{3}\]

OpenStudy (priyar):

i was trying to derive it..

OpenStudy (priyar):

coz one can't remember formulas for each series..right?

OpenStudy (jiteshmeghwal9):

oh sorry i thought that u were trying to solve question

OpenStudy (priyar):

its ok..

OpenStudy (priyar):

adding zero to a series wont make a difference... then the summation part should be same for both (2n+1)^2 and (2n-1)^2..but its not so..why? am i missing something?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & x=1^{2} +3^{2} +5^{2} +\cdots+25^{2}\hspace{.33em}\\~\\ & x+2^{2}+4^{2}+6^{2}+\cdots+24^{2}=1^{2} +2^{2}+3^{2}+4^{2} +5^{2}+ \cdots+25^{2}\hspace{.33em}\\~\\ & x+2^{2}(1+2^{2}+3^{2}+\cdots+12^{2})=1^{2} +2^{2}+3^{2}+4^{2} +5^{2}+ \cdots+25^{2}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (dayakar):

sum of the squares of first natural numbers is\[= \frac{ n(n+1)(2n+1) }{ 6 }\] substitute n=25 u can get the answer

OpenStudy (priyar):

@dayakar these are just odd nos. so we can't do that..

OpenStudy (jiteshmeghwal9):

general term=(\(2n-1)^2\)\[T_n=4n^2-4n+1\]\[\sum4n^2-\sum4n+\sum1\]\[4 \sum n^2-4 \sum n+n\]

OpenStudy (priyar):

ya i sid the same way but with (2n+1)...could u do like that and show me..?

OpenStudy (priyar):

*did

OpenStudy (jiteshmeghwal9):

if u know the sum of first n squares & n natural numbers. u can derive tht

OpenStudy (priyar):

ya..i do know that...but...did u read the abv part...?

OpenStudy (priyar):

adding zero to a series wont make a difference... then the summation part should be same for both (2n+1)^2 and (2n-1)^2..but its not so..why? am i missing something? Also ya... i did the same way but with (2n+1)...could u do like that and show me..?

OpenStudy (priyar):

pls help!

OpenStudy (priyar):

why is it (2k+1)^3 ?

OpenStudy (priyar):

somebody reply pls

OpenStudy (dayakar):

\[4\sum_{?}^{?}n ^{2}-4\sum_{?}^{?}n+n\] \[= \frac{ n(4n ^{2}-1) }{ 3 }\]

OpenStudy (dayakar):

now plug n=25 u can get the answer

OpenStudy (priyar):

what were ur limits for summation?

OpenStudy (dayakar):

in our problem n= 1 to 25

OpenStudy (priyar):

no..its 1 to 13 in the abv. case and anyway i know how to do this problem by this method..

OpenStudy (dayakar):

with out using limits i find general form

OpenStudy (priyar):

my doubt is how to solve this: \[\sum_{0}^{12} (2n + 1)^2\]

OpenStudy (dayakar):

suppose if write n=3 we can get sum of squares of first 3 odd numbers

OpenStudy (dayakar):

expand (2x+1)^2

OpenStudy (priyar):

\[4\sum_{0}^{12} n^2 + 4\sum_{0}^{12} n +1\]

OpenStudy (priyar):

sigma for 1 also

OpenStudy (priyar):

sigma for 1 also

OpenStudy (dayakar):

sigma n^2 = n(n+1)(2n+1)/6 sigma n = n(n+1)/2 sig ma 1 = 12 [ limits are 1 to 12]

OpenStudy (dayakar):

\[4[\sum_{0}^{12}\frac{ n(n+1)(2n+1) }{ 6 } ]+ 4[ \sum_{0}^{12}\frac{ n(n+1) }{ 2 }] + \sum_{0}^{12}1\]

OpenStudy (dayakar):

sorry remove sigma

OpenStudy (priyar):

this summation is only for 1 to n is it the same for 0 to n also?

OpenStudy (dayakar):

addition of 0 to any number makes any difference

OpenStudy (dayakar):

actually i don't know clearly,

OpenStudy (priyar):

yes thats what i am trying to say..then why is the answer diff.?

OpenStudy (dayakar):

what do u get

OpenStudy (dayakar):

can u write the answer

OpenStudy (priyar):

in the method where we used (2n-1)^2 i m getting 2925. in this method i m getting 2924

OpenStudy (dayakar):

i think 2925 is correct

OpenStudy (priyar):

ya but why r we getting to diff answers why is this method wrong?

OpenStudy (dayakar):

in first method limits are 0 to 13 if u write n=0 and n=1 u will get both times 1 so ,we have to subtract 1 from first method answer therefore the answer is 2924

OpenStudy (dayakar):

is it clear

OpenStudy (priyar):

why are the limits 0 to 13 it must be 1 to 13 in the first method..

OpenStudy (priyar):

@dayakar

OpenStudy (priyar):

and the correct answer is...2925

OpenStudy (phi):

The sum \[ \sum_{i=0}^n 1 = n+1 \] not n

OpenStudy (priyar):

yes! at last thank you!!

OpenStudy (priyar):

thanks again @phi .. i was confused for so long...!now i realise where i went wrong!